# Fixed-domain curve counts for blow-ups of projective space

### v3

@@ 2. Case 2
      If instead $f_{*}[C_{{\text{sp}}}]=d^{\prime}\mathsf{H}+k^{\prime}\mathsf{E}^{\vee}$ with $d^{\prime}\geq k^{\prime}\geq 0$ and $d^{\prime}>0$ or $d^{\prime}=k^{\prime}=0$, then Condition (*) follows from Lemma 22.
@@ 2. Step 2
    We verify Condition (**). If some tree $T_{i}$ for $i\in\{1,\ldots,n-m\}$ contains a component $T_{i}^{0}$ whose pushforward via $f$ is not ordinary, then it must be
@@
    which has degree $k^{\prime}(r-1)\geq r-1$. Also, $T_{i}$ has to contain a component $\overline{T_{i}}$ whose pushforward via $f$ is ordinary and which, by Lemma 21, has degree at least $2$. Therefore, $f_{*}[T_{i}]$ has at least degree $r+1$.
$\Diamond$
@@ 2.5. SAE for del Pezzo surfaces
We will use Proposition 19 again. Let $Z\subseteq{\mathcal{M}}_{\Gamma}^{(m,a)}$ be an irreducible component and $[f]\in Z$ a general point.
@@ 1. Step 1
    First, we verify Condition (*). If $m\geq g+1$, then by [17, Proposition 13], we have $h^{1}(C_{{\text{sp}}},f^{*}T_{X})=0$, and so
@@
    If instead $m<g+1$, then $n-m\geq n-g-1$. Observe that, for all $i=1,\ldots,n-m$, the pushforward of some component $\overline{T_{i}}$ of $T_{i}$ must be an ordinary class different from 00. Therefore, by Lemma 21, we have
@@
    where in the second inequality we used (7). In particular, when $g=0$, the right hand side is 00, yielding a contradiction. Therefore, for $g=0$, it must be that $m\geq 1=g+1$ and $h^{1}(C_{{\text{sp}}},f^{*}T_{X})=0$.
@@
    This implies that Condition (*) holds.
@@ 2. Step 2
    We now deal with Condition (**). Let $T_{i}$ be a tree containing a component whose pushforward under $f$ is not ordinary. It still must contain a component $\overline{T_{i}}$ such that $f_{*}([\overline{T_{i}}])$ is an ordinary curve (and so of degree at least $2$ by Lemma 21), and at least one more irreducible component, of degree at least $1$ because $X$ is Fano. Therefore, $\mathrm{deg}(f_{*}([T_{i}])\geq 3=r+1$ and we are done.
@@
2.6. SAE in dimension $3$
Let $\pi:X\rightarrow\mathbb{P}^{3}$ be the blow-up of $\mathbb{P}^{3}$ at $\ell\leq 4$ general points. Write $\mathsf{H},\mathsf{E}_{1},\ldots,\mathsf{E}_{\ell}$ for the divisor classes of the pulled-back hyperplane class and the exceptional divisor, respectively, and write $\mathsf{H}^{\vee},\mathsf{E}^{\vee}_{1},\ldots,\mathsf{E}^{\vee}_{\ell}$ for the corresponding dual basis of $H_{2}(X)$.
In this section, we prove that $X$ satisfies SAE (Theorem 32). Note that such $X$ are not Fano; this is, to our knowledge, the first example of a non-Fano variety whose Tevelev degrees are asymptotically enumerative.
@@ 2.6.1. Low degree effective curves in $\mathbb{P}^{3}$
Lemma 25.
[10, Proposition 4.1] The effective cone of $X$ is linearly generated by $1$-dimensional linear spaces in the exceptional divisors and the strict transforms of $1$-dimensional linear subspaces of $\mathbb{P}^{3}$, possibly passing through the $\ell$ blown up points.
Corollary 26.
@@
Let $\mathbb{T}\subset\mathbb{P}^{3}$ be a torus acting on $\mathbb{P}^{3}$ in such a way that the $q_{j}$ are the torus-fixed points. We also denote by $\mathbb{T}\subset X$ the pullback to $X$, so that $\mathbb{T}$ also acts on $X$. We claim that $C$ lies in the strict transform $\Lambda$ of some torus-fixed plane in $\mathbb{P}^{3}$. Suppose instead that this is not the case. Then, we will show that $\beta$ would be ordinary, and so by Lemma 21, would have anti-canonical degree at least 2, a contradiction.
@@
On the other hand, the degree against $K_{X}^{\vee}$ of the classes $\mathsf{H}^{\vee}+\mathsf{E}_{i}^{\vee}$ and $-\mathsf{E}_{j}^{\vee}$ is $2$, so we find no possibilities for $\beta$. We have reached a contradiction, completing the proof. $\Diamond$
Lemma 29.
@@
and, since $\beta$ is ordinary, by Lemma 20 we have
@@ 2.6.2. Proof of SAE
Assume inequality (7) holds. For integers $a\geq 0$ and $s,m\geq 0$ such that $s+m\leq n$, define
@@
Again, we assume that if $(n-m)+a=0$, that is, if $C$ is irreducible, then (7) is a strict inequality.
Let $Z$ be an irreducible component of ${\mathcal{M}}_{\Gamma}^{(s,m,k)}$. We will prove:
Proposition 31.
If either $n$ is large (depending on $X$ and $g$) or $g=0$ and $n\geq 1$, then the component $Z$ fails to dominate $X^{n}\times\overline{\mathcal{M}}_{g,n}$.
As an immediate corollary, we obtain:
Theorem 32.
@@
We prove Proposition 31 by first making several reductions. Let $[f]\in Z$ be a general point.
Reduction 1: It is enough to prove Proposition 31 for $a=0$.
@@
Because $f_{*}[T_{i}^{\prime}]\cdot K_{X}^{\vee}\geq 0$, this is clear when $m<n$, as we may simply delete the components $T^{\prime}_{i}$, so assume that $m=n$. For the same reason, we can also assume that $a=1$, that $T_{1}^{\prime}$ is irreducible, and that $f_{*}[T_{1}^{\prime}]\cdot K_{X}^{\vee}=0$. It follows from Lemma 28 and stability that
@@
Let $\mathsf{L}$ be the line in $\mathbb{P}^{3}$ through $\pi(\mathsf{E}_{j})$ and $\pi(\mathsf{E}_{j^{\prime}})$, and $\rho:\widetilde{X}\rightarrow X$ be the blow-up of $X$ along the strict transform of $\mathsf{L}$. Call $\overline{\beta}=\beta-f_{*}[T_{1}^{\prime}]$ and let $\overline{Z}\subset{\mathcal{M}}_{g,n}(X,\overline{\beta})$ be the irreducible component to which $[f|_{C_{{\text{sp}}}}]$ belongs. From $\overline{Z}$ we get an irreducible substack
$\widetilde{Z}\subseteq{\mathcal{M}}_{g,n}(\widetilde{X},\widetilde{\beta})$
dominating $\overline{Z}$. Here, $\widetilde{\beta}\in H_{2}(\widetilde{X},\mathbb{Z})$ is an effective curve class with the property that
@@
If $Z$ dominates $X^{n}\times{\mathcal{M}}_{g,n}$, then $\widetilde{Z}$ dominates $\widetilde{X}^{n}\times{\mathcal{M}}_{g,n}$, and applying [17, Proposition 13] twice, we obtain
$\mathrm{dim}(Z)=n+\beta\cdot K_{X}^{\vee}>n+\widetilde{\beta}\cdot K_{\widetilde{X}}^{\vee}\geq\mathrm{dim}(\widetilde{Z}),$
@@
Reduction 2: It is enough to prove Proposition 31 for $s=0$.
@@
If, for some $i\in\{1,\ldots,s\}$, we have $f_{*}[T_{i}]\cdot K_{X}^{\vee}\geq 4=\mathrm{dim}(X)+1$, then arguing as in the proof of Proposition 18 above, we can delete $T_{i}$ from the domain.
@@
be the decomposition in irreducible components of $T_{i}$. Then, by Lemmas 28 and 29, $f_{*}[T_{i}^{j}]$ is a fixed line for all but one $j\in\{0,\ldots.,k_{i}\}$, say except for $j=0$, and $f_{*}[T_{i}^{0}]$ is a pivoting line $\mathsf{H}^{\vee}+\mathsf{E}_{h}^{\vee}$ for some $h\in\{1,\ldots,\ell\}$.
Now, for $[f]\in Z$ such that $f(p_{i})=x_{i}$, we need $x_{i}\in f(T_{i}^{0})$. Thus, the image $f(T_{i}^{0})$ is the strict transform of the unique line between $q_{h}$ and $x_{i}$, while for $j\geq 1$, the reduced image $f(T_{i}^{j})^{\mathrm{red}}$ is either a point or the strict transform of the unique line in $\mathbb{P}^{3}$ through two of the points $q_{1},\ldots,q_{\ell}$. By assumption, there is at least a $j$ for which $[f(T_{i}^{j})^{\mathrm{red}}]$ is a fixed line. However, for general $x_{i}$, two such curves do not meet in $X$. Thus, $Z$ cannot dominate ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$ unless $s=0$. $\Diamond$
@@
If $f_{*}[T_{i}]\cdot K_{X}^{\vee}\geq 4$, then arguing as in the proof of Proposition 18, we can delete $T_{i}$. By Lemma 21, we can assume that if (11) is the decomposition in irreducible components of $T_{i}$, then $f(T_{i}^{j})$ is a point unless $j=0$, and $f(T_{i}^{0})$ is the strict transform of a line through one of the points $q_{1},\ldots,q_{\ell}$. In particular, replacing $T_{i}$ with $T_{i}^{0}$, it is enough to prove Proposition 31 when $T_{i}=T_{i}^{0}$. $\Diamond$
@@
Reduction 4: It is enough to prove the following analogue of condition (*):
@@
    The following inequality holds:
    $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}$ $]))-n$
    $\leq-(g+1);$
@@
The proof is the same as that of Proposition 17. $\Diamond$
Proof of Proposition 32.
It is enough to prove that condition (*’) holds whenever $n$ is large (depending on $X$ and $g$) and whenever $g=0$ and $n\geq 1$. We distinguish two cases.
@@ 1. Case 1
    and so by Equation (1),
    $3(n+g-1)=\beta\cdot K_{X}^{\vee}=4n-2(n-k^{\prime})=2n+2k^{\prime},$
@@
    $k^{\prime}=\frac{n}{2}+\frac{3}{2}(g-1).$
    Proceeding as in Theorem 23, we find
    $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))$ $\leq 2g-1+k^{\prime}$
    $\leq\frac{7}{2}g-\frac{5}{2}+\frac{n}{2}$
    which implies the required inequality (*’).

### v4

@@ 2. Case 2
      If instead $f_{*}[C_{{\text{sp}}}]=d^{\prime}\mathsf{H}^{\vee}+k^{\prime}\mathsf{E}^{\vee}$ with $d^{\prime}\geq k^{\prime}\geq 0$ and $d^{\prime}>0$ or $d^{\prime}=k^{\prime}=0$, then Lemma 2.7 gives that $h^{1}(C_{{\text{sp}}},f^{*}T_{X})\leq rg$. Therefore,
      $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))\leq h^{1}(C_{{\text{sp}}},f^{*}T_{X})\leq n-(g+1).$
      as long as $n\geq rg+g+1$.
    Combining cases 1 and 2, we see that condition (*) holds for
    $n\geq C^{\prime}(X,g):=\max((r^{2}+3r+1)g,rg+g+1)$
    for all $g$, and in particular for $n\geq 3$ if $g=0$.
@@ 2. Step 2
    We verify condition (**). Here, we require no assumption on $n$. If some tree $T_{i}$ for $i\in\{1,\ldots,n-m\}$ contains a component $T_{i}^{0}$ whose pushforward via $f$ is not ordinary, then it must be
@@
    which has degree $k^{\prime}(r-1)\geq r-1$. Also, $T_{i}$ has to contain a component $\overline{T_{i}}$ whose pushforward via $f$ is ordinary and which, by Lemma 2.6, has degree at least $2$. Therefore, $f_{*}[T_{i}]$ has at least degree $r+1$.
Therefore the criterion of Proposition 2.4 holds, and $X$ satisfies SAE. $\Diamond$
@@ 2.5. SAE for del Pezzo surfaces
We adopt the setup of Proposition 2.4 again.
@@ 1. Step 1
    First, we consider condition (*). If $m\geq g+1$, then by [16, Proposition 13], we have $h^{1}(C_{{\text{sp}}},f^{*}T_{X})=0$, and so
@@
    If instead $m<g+1$, then nm>ng1n-m>n-g-1. Observe that, for all $i=1,\ldots,n-m$, the pushforward of some component $\overline{T_{i}}$ of $T_{i}$ must be an ordinary class different from 0. Therefore, by Lemma 2.6, we have
@@
    where in the second inequality we used (8). In particular, when $g=0$, the right hand side is 0, yielding a contradiction. Therefore, for $g=0$, it must be that $m\geq 1=g+1$ and $h^{1}(C_{{\text{sp}}},f^{*}T_{X})=0$.
@@
    This implies that condition (*) holds, for all $n$ sufficiently large (when $g$ is arbitrary), and for all $n\geq 3$ (when $g=0$).
@@ 2. Step 2
    We now deal with condition (**). Let $T_{i}$ be a tree containing a component whose pushforward under $f$ is not ordinary. It still must contain a component $\overline{T_{i}}$ such that $f_{*}([\overline{T_{i}}])$ is an ordinary curve (and so of degree at least $2$ by Lemma 2.6), and at least one more irreducible component, of degree at least $1$ because $X$ is Fano. Therefore, $\mathrm{deg}(f_{*}([T_{i}])\geq 3=r+1$ and we are done.
@@
2.6. SAE in dimension 33
Let $\pi:X\rightarrow\mathbb{P}^{3}$ be the blow-up of $\mathbb{P}^{3}$ at $\ell\leq 4$ general points. Write $\mathsf{H},\mathsf{E}_{1},\ldots,\mathsf{E}_{\ell}$ for the divisor classes of the pulled-back hyperplane class and the exceptional divisors, respectively, and write $\mathsf{H}^{\vee},\mathsf{E}^{\vee}_{1},\ldots,\mathsf{E}^{\vee}_{\ell}$ for the corresponding dual basis of $H_{2}(X)$.
In this section, we prove that $X$ satisfies SAE (Theorem 2.17). Note that such $X$ are not Fano; this is, to our knowledge, the first example of a non-Fano variety whose Tevelev degrees are asymptotically enumerative.
@@ 2.6.1. Low degree effective curves in $\mathbb{P}^{3}$
Lemma 2.10.
[11, Proposition 4.1] The effective cone of $X$ is linearly generated by $1$-dimensional linear spaces in the exceptional divisors and the strict transforms of $1$-dimensional linear subspaces of $\mathbb{P}^{3}$, possibly passing through the $\ell$ blown up points.
Corollary 2.11.
@@
Let $\mathbb{T}\subset\mathbb{P}^{3}$ be a torus acting on $\mathbb{P}^{3}$ in such a way that the $q_{j}$ are the torus-fixed points. We also denote by $\mathbb{T}\subset X$ the pullback to $X$, so that $\mathbb{T}$ also acts on $X$. We claim that $C$ lies in the strict transform $\Lambda$ of some torus-fixed plane in $\mathbb{P}^{3}$. Suppose instead that this is not the case. Then, we will show that $\beta$ would be ordinary, and so by Lemma 2.6, would have anti-canonical degree at least 2, a contradiction.
@@
On the other hand, the degree against $K_{X}^{\vee}$ of the classes $\mathsf{H}^{\vee}+\mathsf{E}_{i}^{\vee}$ and $\mathsf{E}_{j}^{\vee}$ is $2$, so we find no possibilities for $\beta$. We have reached a contradiction, completing the proof. $\Diamond$
Lemma 2.14.
@@
and, since $\beta$ is ordinary, by Lemma 2.5 we have
@@ 2.6.2. Proof of SAE
For integers $a\geq 0$ and $s,m\geq 0$ such that $s+m\leq n$, define
@@
Again, we assume that if (nm)+a=0(n-m)+a=0, that is, if $C$ is irreducible, then (8) is a strict inequality.
Let $Z$ be an irreducible component of ${\mathcal{M}}_{\Gamma}^{(s,m,k)}$. We will prove the following.
Proposition 2.16.
Assume that, for all $g\geq 0$, there exists an integer $C^{\prime}(X,g)$ with the following property. Suppose that $n,\beta$ satisfy $n\geq C^{\prime}(X,g)$ and that the inequality
(14) $\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))\leq\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n})$
holds. For any $s,m,k$, let $Z\subset{\mathcal{M}}_{\Gamma}^{(s,m,k)}\subset\overline{\mathcal{M}}_{g,n}(X,\beta)$ be an irreducible component. Suppose further that if $(m,a)=(n,0)$, then (14) is a strict inequality. Then, $Z$ fails to dominate $\overline{\mathcal{M}}_{g,n}\times X^{n}$.
The inequality (14) is equivalent to $\beta\cdot K_{X}^{\vee}\leq 3(n+g-1)$. As an immediate corollary to Proposition 2.16, we obtain the following.
Theorem 2.17.
@@
We prove Proposition 2.16 by first making several reductions. Let $f]\in Z$ be a general point.
Reduction 1: It is enough to prove Proposition 2.16 for $a=0$.
@@
Because $f_{*}[T_{i}^{\prime}]\cdot K_{X}^{\vee}\geq 0$, this is clear when $m<n$, as we may simply delete the components $T^{\prime}_{i}$, so assume that $m=n$. For the same reason, we can also assume that $a=1$, that $T_{1}^{\prime}$ is irreducible, and that $f_{*}[T_{1}^{\prime}]\cdot K_{X}^{\vee}=0$. It follows from Lemma 2.13 and stability that
@@
Let $\mathsf{L}$ be the line in $\mathbb{P}^{3}$ through $\pi(\mathsf{E}_{j})$ and $\pi(\mathsf{E}_{j^{\prime}})$, and $\rho:\widetilde{X}\rightarrow X$ be the blow-up of $X$ along the strict transform of $\mathsf{L}$. Call $\overline{\beta}=\beta-f_{*}[T_{1}^{\prime}]$ and let $\overline{Z}\subset{\mathcal{M}}_{g,n}(X,\overline{\beta})$ be the irreducible component to which $f|_{C_{{\text{sp}}}}]$ belongs. From $\overline{Z}$ we get an irreducible substack
[…]
@@
If $Z$ dominates $X^{n}\times{\mathcal{M}}_{g,n}$, then $\widetilde{Z}$ dominates $\widetilde{X}^{n}\times{\mathcal{M}}_{g,n}$, and applying [16, Proposition 13] twice, we obtain
[…]
@@
Reduction 2: It is enough to prove Proposition 2.16 for $s=0$.
@@
If, for some $i\in\{1,\ldots,s\}$, we have $f_{*}[T_{i}]\cdot K_{X}^{\vee}\geq 4=\mathrm{dim}(X)+1$, then arguing as in the proof of Proposition 2.4 above, we can delete $T_{i}$ from the domain.
@@
be the decomposition in irreducible components of $T_{i}$. Then, by Lemmas 2.13 and 2.14, $f_{*}[T_{i}^{j}]$ is a fixed line for all but one $j\in\{0,\ldots.,k_{i}\}$, say except for $j=0$, and $f_{*}[T_{i}^{0}]$ is a pivoting line $\mathsf{H}^{\vee}+\mathsf{E}_{h}^{\vee}$ for some $h\in\{1,\ldots,\ell\}$.
Now, for $f]\in Z$ such that $f(p_{i})=x_{i}$, we need $x_{i}\in f(T_{i}^{0})$. Thus, the image $f(T_{i}^{0})$ is the strict transform of the unique line between $q_{h}$ and $x_{i}$, while for $j\geq 1$, the reduced image $f(T_{i}^{j})^{\mathrm{red}}$ is either a point or the strict transform of the unique line in $\mathbb{P}^{3}$ through two of the points $q_{1},\ldots,q_{\ell}$. By assumption, there is at least a $j$ for which $f(T_{i}^{j})^{\mathrm{red}}]$ is a fixed line. However, for general $x_{i}$, two such curves do not meet in $X$. Thus, $Z$ cannot dominate $\overline{\mathcal{M}}_{g,n}\times X^{n}$ unless $s=0$. $\Diamond$
@@
If $f_{*}[T_{i}]\cdot K_{X}^{\vee}\geq 4$, then arguing as in the proof of Proposition 2.4, we can delete $T_{i}$. By Lemma 2.6, we can assume that if (15) is the decomposition in irreducible components of $T_{i}$, then $f(T_{i}^{j})$ is a point unless $j=0$, and $f(T_{i}^{0})$ is the strict transform of a line through one of the points $q_{1},\ldots,q_{\ell}$. In particular, replacing $T_{i}$ with $T_{i}^{0}$, it is enough to prove Proposition 2.16 when $T_{i}=T_{i}^{0}$. $\Diamond$
@@
Reduction 4: It is enough to prove the following analog of condition (*).
@@
    The following inequality holds.
    $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))\leq n-(g+1).$
@@
The proof is the same as that of Proposition 2.3. $\Diamond$
Proof of Proposition 2.16.
It is enough to prove that condition (*’) holds whenever $n$ is large (depending on $X$ and $g$) and whenever $g=0$ and $n\geq 3$. There are two cases.
@@ 1. Case 1
    and so by (14),
    $3(n+g-1)\geq\beta\cdot K_{X}^{\vee}=4n-2(n-k^{\prime})=2n+2k^{\prime},$
@@
    $k^{\prime}\leq\frac{n}{2}+\frac{3}{2}(g-1).$
    Proceeding as in Theorem 2.8, we find
    $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))$
    $\leq$ $\,\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{2},k^{\prime}))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))+2g$
    $=$ $\,[3(k^{\prime}-g+1)-1]-[4(-g+1)+2k^{\prime}-1]+2g$
    $=$  $\,3g-1+k^{\prime}$
    $\leq$ $\,\frac{9}{2}g-\frac{5}{2}+\frac{1}{2}n$
    which implies the required inequality (*’) if $n\geq 11g-7$ or if $g=0$ and $n\geq 3$.